A plant using process costing has 4,000 units in beginning work-in-process, and started 15,000 more. If 5,000 units are in the ending work-in-process then: For Answer Scoring 1. How many units are there to account for? 1. 2. How many units are transferred using weighted average method? 2. 3. How many units are transferred using first-in, first out method? 3. 4. How many units were both started and completed during this period? 4.
Answer #1

1. Units to account for = Units in beginning work in process + Units started

= 4,000 + 15,000

= 19,000

2. Units transferred using weighted average method = Units that are completed + Units in ending work in process

= (19,000 - 5,000) + 5,000

= 19,000

3. Units transferred using first in first out = Units in beginning work in process + Units started and completed + Units in ending work in process

= 4,000 + (19,000 - 5,000 - 4,000) + 5,000

= 19,000

4. Units stared and completed = Units to account for - Units in ending work in process - Units in beginning work in process

= 19,000 - 5,000 - 4,000

= 10,000